OpenStudy (anonymous):

Rotate the axes to eliminate the xy-term in the equation. Then write the equation in standard form. 97x^2 + 130xy + 97y^2 - 2592 = 0

OpenStudy (anonymous):

97x² + 130xy + 97y² - 2592 = 0 cot(2θ) = (97-97)/130 ... ⇒ θ = π/4 x = X cos(θ) - Y sin(θ) and y = X sin(θ) + Y cos(θ) 97/2 (X - Y)² + 65 (X - Y)(X + Y) + 97/2 (X + Y)² - 2592 = 0 162 X² + 32 Y² = 2592 X²/16 + Y²/81 = 1 Answer: X²/16 + Y²/81 = 1

OpenStudy (anonymous):

@ziko1995 can you help me with these too? x2 + 10xy + y2 - 8 = 0 145x2 - 34xy + 145y2 - 10368 = 0

OpenStudy (anonymous):

x² + 10xy + y² - 8 = 0 cot(2θ) = (1-1)/10 ... ⇒ θ = π/4 1/2 (X - Y)² + 5 (X - Y)(X + Y) + 1/2 (X + Y)² - 8 = 0 3 X² - 2 Y² = 4 X²/(4/3) - Y²/2 = 1 Answer: X²/(4/3) - Y²/2 = 1 145x² - 34xy + 145y² - 10368 = 0 cot(2θ) = (145-145)/-34 ... ⇒ θ = π/4 145/2 (X - Y)² - 17 (X - Y)(X + Y) + 145/2 (X + Y)² - 10368 = 0 128 X² + 162 Y² = 10368 X²/81 + Y²/64 = 1 Answer: X²/81 + Y²/64 = 1

OpenStudy (anonymous):

wow thanks!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

If u wanna be friends ; My facebook account : http://www.facebook.com/zakaria.moufrad.8

OpenStudy (anonymous):

alright, i cant go on facebook, twitter, etc from this computer but i will later